# Reinforcement Learning from Probabilistic Forecasts for Safe Decision-Making via Conditional Value-at-Risk Planning

### v1

@@ 5.1 Point forecast accuracy
Table 3: S&P 500 — RMSE (mean $\pm$ SD).
Model 1 5 30
@@
Table 4: S&P 500 — MAE (mean $\pm$ SD).
Model 1 5 30
@@
Table 5: S&P 500 — sMAPE (mean $\pm$ SD).
Model 1 5 30
@@
Table 6: H&M — RMSE (mean $\pm$ SD).
Model 1 7 28
@@
Table 7: H&M — MAE (mean $\pm$ SD).
Model 1 7 28
@@
Table 8: H&M — sMAPE (mean $\pm$ SD).
Model 1 7 28
@@ 5.2 Probabilistic forecast evaluation
The UAMDP delivers marked improvements in probabilistic forecast quality compared to baselines. For the S&P 500 dataset, it reduces CRPS by 29% relative to the best-performing baseline (TFT); for H&M retail demand, the CRPS reduction is 34%. In terms of empirical coverage, the UAMDP consistently aligns with nominal levels, achieving 80.8% for the 80% interval and 95.6% for the 95% interval. By contrast, Prophet systematically under-covers in the upper tail, reaching only 73% coverage at the 80% level.
@@ 5.3 Reinforcement-learning performance
Table 11 summarizes performance across key financial metrics, with UAMDP achieving a Sharpe ratio of 1.74, marking a 13.2% improvement over the best-performing baseline (PatchTST, Sharpe = 1.54). Maximum drawdown is simultaneously reduced from 12.1% to 5.1%, indicating a substantially improved risk–return profile. These differences are statistically significant based on Wilcoxon signed-rank tests ($p<0.01$ for all pairwise comparisons).
@@ 5.4 Risk-sensitive decision analysis
The trade-off between mean reward and CVaR across different risk thresholds ($\alpha\in\{0.01,0.05,0.10\}$) is illustrated in Fig. 6. UAMDP consistently outperforms baselines by achieving higher mean rewards at comparable or lower CVaR levels. At the $\alpha=0.05$ risk level with $\lambda=0.8$, UAMDP attains a CVaR of $0.84$, significantly outperforming TFT’s $0.67$ (a 26% improvement). The heatmap in Fig. 6 reveals a broad diagonal region of near-optimal performance, indicating model robustness and low sensitivity to hyperparameter selection.
@@ 5.5 Ablation study
Table 13 presents the impact on performance metrics of removing key model components. Eliminating Thompson sampling significantly reduces the service level by 0.067 and the Sharpe ratio by 0.262, emphasizing the importance of structured exploration, even within seemingly stationary series. Removing Bayesian belief filtering results in the most substantial degradation in Sharpe (0.299-0.299), reinforcing the critical role of accurate posterior tracking. Finally, dropping the CVaR objective reduces tail-risk performance (CVaR drops from 0.84 to 0.63) but negligibly affects mean reward, clarifying the CVaR objective’s role as a mechanism primarily for shaping risk rather than directly driving profits.
@@ 6 Discussion
In this study, we set out to evaluate whether the UAMDP framework can materially enhance both forecast fidelity and downstream control performance in volatile, high-stakes environments. We hypothesized that by embedding a Bayesian time-series forecasting module within a posterior-sampling RL loop and layering a CVaR objective on top, the framework’s agents could achieve three objectives concurrently-reduce prediction error, improve economic outcomes, and mitigate tail risk—all without incurring a penalty in sample efficiency. Empirical evidence from two disparate domains, high-frequency equity trading and low-frequency inventory control, supports this hypothesis, highlighting both the generality and effectiveness of the proposed approach. The principal findings are summarized below. The following are the main findings:
@@
    Posterior sampling enables risk-sensitive exploration. Ablation of the Thompson sampling component led to diminished service-level performance and a drop in the trading Sharpe ratio (Table 7), reaffirming that posterior-guided exploration remains valuable even in quasi-stationary environments. This aligns with theoretical regret bounds from the literature on posterior sampling for reinforcement learning (PSRL) and extends their validity to million-step, partially observable tasks with structured priors.
@@
    CVaR planning constrains adverse tail events. Under fixed mean-return levels, the UAMDP achieved a $\mathrm{CVaR}_{0.05}$ of $0.84$, as compared to $0.67$ for TFT, representing a 26% reduction in extreme losses, while maintaining feasibility under domain-specific constraints (Fig. 6). A heat map over $(\alpha,\lambda)$ pairs revealed a broad diagonal ridge of near optimality, suggesting robustness to moderate hyperparameter misspecification.
@@
A principal limitation of the present framework lies in the use of approximate inference and truncated planning. Real-time constraints necessitated using a 256-particle filter and capping Monte Carlo tree search rollouts at 128. Although Section 3.2.3 bounds the resulting error at $\mathcal{O}(\varepsilon_{p}+\varepsilon_{f})$, degradation in long-horizon calibration was evident beyond $H>64$, suggesting underestimation of posterior variance in highly non-stationary or heavy-tailed environments.
@@
Methodologically, extending the UAMDP to multi-agent, partially observable environments—including competitive trading, cooperative routing, and shared-resource scheduling—would test its scalability under strategic interactions. Integration with observation-efficient offline RL may also enhance safety by leveraging logged trajectories. Finally, replacing static CVaR with dynamic distortion or spectral risk measures offers a path toward fine-grained, time-consistent tail control, particularly in sectors with rolling-reserve or capacity-constrained obligations.
@@
Proof. Let $t\geq 0$ and let $a_{t}\sim\pi(\,\cdot\mid H_{t})$, where $H_{t}=(o_{t},a_{t-1})$ is the observable history. Let $\mathbf{1}_{B}$ be the indicator of a measurable set $B$, defined by
$\mathbf{1}_{B}\bigl(b_{t},a_{t},o_{t+1}\bigr)=\begin{cases}1,&\text{if }\mathrm{Bayes}(b_{t},a_{t},o_{t+1})\in B,\\ 0,&\text{otherwise.}\end{cases}$
Since the latent parameter $\theta$ is drawn once at $t=0$ and remains constant, the predictive density of the next observation decomposes as
$p(o_{t+1}\mid H_{t},a_{t})=\int_{\theta}p(o_{t+1}\mid\theta,H_{t},a_{t})\,\beta_{t}(\theta)\,d\theta.$
Exact filtering (Assumption A2) updates the belief via Bayes’ rule:
$\beta_{t+1}(\theta)=\frac{p(o_{t+1}\mid\theta,H_{t},a_{t})\,\beta_{t}(\theta)}{\displaystyle\int_{\theta^{\prime}}p(o_{t+1}\mid\theta^{\prime},H_{t},a_{t})\,\beta_{t}(\theta^{\prime})\,d\theta^{\prime}}.$
Hence, for any set $B$, it follows that
$\Pr\!\bigl\{\beta_{t+1}\in B\mid H_{t},a_{t}\bigr\}=\int_{o_{t+1}}\mathbf{1}_{B}\!\bigl(b_{t},a_{t},o_{t+1}\bigr)\,p(o_{t+1}\mid H_{t},a_{t})\,do_{t+1}.$
The integrand depends on the history only through $(\beta_{t},a_{t})$; therefore
$\Pr\!\bigl\{\beta_{t+1}\in B\mid H_{t},a_{t}\bigr\}=\Pr\!\bigl\{\beta_{t+1}\in B\mid\beta_{t},a_{t}\bigr\},$
proving the Markov property with kernel $\tilde{p}$. $\square$
@@
Theorem 1 (Bayes-Optimality under Exact Inference). Let $b_{0}$ be the prior over latent parameters $\theta$, and let $s_{0}\in\mathcal{S}$ be any initial state. Then the policy executed by UAMDP, $\pi^{U}$, achieves the Bayes value, so its cumulative Bayes regret is exactly zero:
@@
Using the Bayes value equality inside the expectation,
@@
Moreover, since in each episode $G_{k}=V^{\pi^{*}_{M(\theta_{k})}}_{M(\theta_{k})}(s_{0})$ deterministically, the summands are equal to zero episode-wise, so $\mathrm{regret}_{K}=0$. $\square$
Theorem 2 (Finite-Time Regret under Approximate Inference). Let $\pi^{U}$ be the policy produced by UAMDP. Under A1–A3 (realizability, particle filter with $N$ particles, and depth-$L$ Monte-Carlo tree search), for any $\delta\in(0,1)$, with probability at least $1-\delta$:
@@
where $C_{1},C_{2}$ are absolute constants depending on $\gamma$ and $H$.
Proof. Let $M(\theta)=(\mathcal{S},\mathcal{A},P_{\theta},R,\gamma)$ be the true MDP for the latent parameter $\theta^{\star}$. Denote by $b_{t}$ the exact Bayes belief, and by $\hat{b}_{t}$ the belief maintained by the particle filter with $N$ particles. State distributions induced by the optimal and implemented policies are $d_{t}^{\star}$ and $d_{t}^{U}$, respectively. Let $\mathrm{regret}_{H}$ be the cumulative Bayes regret defined by
@@
We decompose the regret into [40]:
@@
By Theorem 1, with probability at least $1-\tfrac{\delta}{3}$, it follows that
$R_{\mathrm{ps}}\;\leq\;\sqrt{HSA\,\ln\!\frac{3}{\delta}}.$
By [41], we obtain
$R_{\mathrm{pf}}\;\leq\;\frac{C_{1}}{\sqrt{N}},\qquad C_{1}=\frac{\gamma^{H}}{(1-\gamma)^{2}}.$
Since Assumption A2 states that the depth-$L$ planner returns an $\mathcal{O}(1/L)$-optimal value estimate, it holds that
$R_{\mathrm{pl}}\;\leq\;\frac{C_{2}}{L},\qquad C_{2}=H.$
Taking a union bound over these events yields probability $1-\delta$ that all three bounds hold simultaneously.
@@
CRediT authorship contribution statement
  • Michal Koren: data curation, methodology, experiment, writing, editing.
  • Or Peretz: data curation, methodology, experiment, visualization, writing, editing.
  • Tai Dinh: visualization, writing, editing, validation
  • Philip S. Yu: writing, editing, validation

### v2

@@ 5.1 Point forecast accuracy
Table 3: S&P 500 — RMSE (mean $\pm$ SD) at 1-, 5-, and 30-minute horizons.
Model 1-min 5-min 30-min
@@
Table 4: S&P 500 — MAE (mean $\pm$ SD) at 1-, 5-, and 30-minute horizons.
Model 1-min 5-min 30-min
@@
Table 5: S&P 500 — sMAPE (mean $\pm$ SD, %) at 1-, 5-, and 30-minute horizons.
Model 1-min 5-min 30-min
@@
Table 6: H&M — RMSE (mean $\pm$ SD) at 1-, 7-, and 28-day horizons.
Model 1-day 7-day 28-day
@@
Table 7: H&M — MAE (mean $\pm$ SD) at 1-, 7-, and 28-day horizons.
Model 1-day 7-day 28-day
@@
Table 8: H&M — sMAPE (mean $\pm$ SD, %) at 1-, 7-, and 28-day horizons.
Model 1-day 7-day 28-day
@@ 5.2 Probabilistic forecast evaluation
The UAMDP delivers marked improvements in probabilistic forecast quality compared to baselines. For the S&P 500 dataset, it reduces CRPS by 29% relative to the best-performing baseline (TFT); for H&M retail demand, the CRPS reduction is 34%. In terms of empirical coverage, the UAMDP aligns closely with the nominal targets. On the S&P 500 dataset it attains 80.8% versus a nominal 80% and 90.2% versus 95% coverage, while on the H&M dataset the corresponding values are 76.8% versus 80% and 93.2% versus 95% (Tables 9–10). By contrast, Prophet systematically under-covers in the upper tail, reaching only 77.9% coverage at the 80% level on H&M (Table 9).
@@ 5.3 Reinforcement-learning performance
Table 11 summarizes performance across key financial metrics. UAMDP achieves a mean daily Sharpe ratio of 1.74, as defined in Section 4.3, representing a 13.2% improvement over the best-performing baseline (PatchTST, daily Sharpe = 1.54). Maximum drawdown is simultaneously reduced from 12.1% to 5.1%, and the increase in turnover (from 1.92 to 3.01) reflects more active rebalancing rather than excessive churning, as it remains within the transaction-cost regime specified in the environment. These differences are statistically significant based on Wilcoxon signed-rank tests ($p<0.01$ for all pairwise comparisons).
@@ 5.4 Risk-sensitive decision analysis
The trade-off between mean reward and CVaR across different risk thresholds ($\alpha\in\{0.01,0.05,0.10\}$) is illustrated in Fig. 6. UAMDP consistently outperforms baselines by achieving higher mean rewards at comparable or lower CVaR levels. At the $\alpha=0.05$ risk level with $\eta=0.8$, UAMDP attains a CVaR of $0.84$, significantly outperforming TFT’s $0.67$ (a 26% improvement). The heatmap in Fig. 6 reveals a broad diagonal region of near-optimal performance, indicating model robustness and low sensitivity to hyperparameter selection.
@@ 5.5 Ablation study
Table 13 presents the impact on performance metrics of removing key model components. Eliminating Thompson sampling significantly reduces the service level by 0.067 and the Sharpe ratio by 0.262, emphasizing the importance of structured exploration even within seemingly stationary series. Removing Bayesian belief filtering results in the most substantial degradation in Sharpe (0.299-0.299), reinforcing the critical role of accurate posterior tracking. Dropping the CVaR objective leads to a similar reduction in Sharpe and service level as the Thompson-sampling ablation, but with a pronounced worsening in tail-risk performance (the CVaR of returns drops from 0.84 to 0.63). This pattern suggests that, in our setting, exploration and risk-sensitive evaluation both contribute to the overall risk–return profile in a coupled way: CVaR primarily shapes the distribution of outcomes and downside protection, while Thompson sampling affects where in state space that risk is taken.
Our ablation analysis focuses on the internal components of the UAMDP stack rather than on an exhaustive comparison with alternative risk-aware RL algorithms. In particular, we do not benchmark against methods that combine different CVaR-style objectives with their own bespoke forecasting or value-function architectures. Designing fully comparable baselines that incorporate calibrated probabilistic forecasts, posterior-driven exploration, and coherent risk measures in a unified way would require substantial additional engineering and is beyond the scope of this work. We therefore regard Table 13 as providing qualitative insight into the contributions of Thompson sampling, Bayesian belief filtering, and CVaR evaluation within UAMDP, and leave a broader empirical study of risk-sensitive RL alternatives to future work.
For brevity, Table 13 reports only Sharpe ratio (S&P 500) and service level (H&M) as summary indicators of risk-adjusted performance and service quality. A more comprehensive ablation across additional metrics such as maximum drawdown, turnover, and cost components would provide a finer-grained view of each module’s contribution and is an important direction for future empirical work.
@@ 6 Discussion
In this study, we set out to evaluate whether the UAMDP framework can materially enhance both forecast fidelity and downstream control performance in volatile, high-stakes environments. We hypothesized that by embedding a Bayesian time-series forecasting module within a posterior-sampling RL loop and layering a CVaR objective on top, the framework’s agents could concurrently achieve three objectives—reduce prediction error, improve economic outcomes, and mitigate tail risk—without incurring a penalty in sample efficiency. Empirical evidence from two disparate domains, high-frequency equity trading and low-frequency inventory control, supports this hypothesis, highlighting both the generality and effectiveness of the proposed approach. The main findings are summarized below.
@@
    Posterior sampling enables risk-sensitive exploration. Ablation of the Thompson sampling component led to diminished service-level performance and a drop in the trading Sharpe ratio (Table 7), reaffirming that posterior-guided exploration remains valuable even in quasi-stationary environments. This aligns with theoretical regret bounds from the literature on posterior sampling for reinforcement learning (PSRL) and extends their qualitative insights to long-horizon, partially observable tasks with structured priors.
@@
    CVaR planning constrains adverse tail events. Under fixed mean-return levels, the UAMDP achieved a $\mathrm{CVaR}_{0.05}$ of $0.84$, as compared to $0.67$ for TFT, representing a 26% reduction in extreme losses, while maintaining feasibility under domain-specific constraints (Fig. 6). A heat map over $(\alpha,\eta)$ pairs revealed a broad diagonal ridge of near optimality, suggesting robustness to moderate hyperparameter misspecification.
@@
A principal limitation of the present framework lies in the use of approximate inference and truncated planning. Real-time constraints necessitated using a 256-particle filter and capping Monte Carlo tree search rollouts at 128. Although Section 3.2.3 bounds the resulting error at $\mathcal{O}(\varepsilon_{p}+\varepsilon_{f})$, our empirical results indicate that very large planning horizons can still be problematic in practice: when $H$ is pushed well beyond the range used in our main experiments (e.g., $H>64$), overall decision quality deteriorates and the forecasts become visibly overconfident relative to realized outcomes. This behavior is consistent with a loss of long-horizon calibration arising from approximate filtering and model misspecification, and highlights that Assumption A2 (exact filtering) should be interpreted as an idealization rather than as a literal description of the deployed system. A more systematic analysis of how calibration varies with the planning horizon is left to future work.
@@
Methodologically, extending the UAMDP to multi-agent, partially observable environments, including competitive trading, cooperative routing, and shared-resource scheduling, would test its scalability under strategic interactions. Integration with observation-efficient offline RL may also enhance safety by leveraging logged trajectories. Finally, replacing static CVaR with dynamic distortion or spectral risk measures offers a path toward fine-grained, time-consistent tail control, particularly in sectors with rolling-reserve or capacity-constrained obligations.
7 Conclusion
This paper has explored how calibrated probabilistic forecasts can be coupled with risk-aware reinforcement learning to support safer, more robust sequential decision-making. The proposed UAMDP framework treats the forecaster as a modular likelihood model, augments the state with Bayesian beliefs, and exposes a CVaR-based risk layer to shape tail outcomes, yielding a unified view of forecasting and control.
Across equity trading and retail inventory domains, our instantiations of UAMDP with Gaussian-process and transformer-based forecasters and a Monte Carlo tree search planner improved risk-sensitive performance relative to strong forecasting and control baselines, while respecting domain-specific constraints. These results suggest that bringing forecast calibration, posterior-driven exploration, and coherent risk measures into a single closed-loop architecture is a promising direction for deploying uncertainty-aware machine learning in operational settings.
@@ 7 Conclusion
Proof. Fix $t\geq 0$. Under policy $\pi$, the action is selected as $a_{t}\sim\pi(\cdot\mid b_{t},s_{t})$, so it depends on the past only through the current hyper-state $(b_{t},s_{t})$. Given $(b_{t},s_{t},a_{t})$, the next observation $x_{t+1}$ has predictive density
$p(x_{t+1}\mid b_{t},s_{t},a_{t})=\int_{\Theta}p_{\phi}(x_{t+1}\mid s_{t},a_{t},\theta)\,b_{t}(\theta)\,d\theta.$
The belief and state are then updated deterministically by
$b_{t+1}=B(b_{t},s_{t},a_{t},x_{t+1}),$
here $B$ and $\tau$ are as defined in the main text.. Therefore, for any measurable set $C\subseteq\mathcal{B}\times\mathcal{S}$, the conditional distribution of $(b_{t+1},s_{t+1})$ given the entire history depends only on $(b_{t},s_{t})$ (and $a_{t}$), which proves that $\{(b_{t},s_{t})\}_{t\geq 0}$ is Markov with transition kernel $\tilde{p}$. $\square$
@@
Theorem 1 (Bayes-optimality under exact inference). Let $b_{0}$ be the prior over latent parameters $\theta$, and let $s_{0}\in\mathcal{S}$ be any initial state. Then the policy executed by UAMDP, denoted $\pi^{U}$, is Bayes-optimal: its Bayes value coincides with the Bayes-optimal value,
@@
In particular, the expected cumulative Bayes regret of $\pi^{U}$ is zero.
@@
Using the Bayes value equality inside the expectation, we obtain
@@
Thus $\pi^{U}$ attains the Bayes value and has zero Bayes regret in expectation. $\square$
Theorem 2 (Finite-time regret under approximate inference). Let $\pi^{U}$ be the policy produced by UAMDP. Under A1–A3 (realizability, a particle filter with $N$ particles, and a depth-$L$ Monte Carlo tree search planner), for any $\delta\in(0,1)$, with probability at least $1-\delta$:
@@
where $C_{1},C_{2}>0$ are absolute constants depending only on $\gamma$ and $H$.
Proof. Let $\mathcal{M}(\theta)=(\mathcal{S},\mathcal{A},P_{\theta},R,\gamma)$ denote the MDP parameterized by $\theta$, and let $\theta^{\star}$ be the true latent parameter. Denote by $b_{t}$ the exact Bayes belief, and by $\hat{b}_{t}$ the belief maintained by the particle filter with $N$ particles. State distributions induced by the optimal and implemented policies are $d_{t}^{\star}$ and $d_{t}^{U}$, respectively. Let $\mathrm{regret}_{H}$ be the cumulative Bayes regret defined by
@@
We decompose the regret as in [russo2014learning]:
@@
By the posterior-sampling regret analysis of [russo2014learning] and the bounded-reward Assumption A5, there exists a universal constant $C>0$ such that, with probability at least $1-\tfrac{\delta}{3}$,
$R_{\mathrm{ps}}\;\leq\;C\,\sqrt{HSA\,\ln\!\frac{3}{\delta}}.$
By standard Feynman–Kac particle-filter concentration bounds [moral2004feynman], there exists $C_{1}>0$ depending only on $(\gamma,H)$ such that
$R_{\mathrm{pf}}\;\leq\;\frac{C_{1}}{\sqrt{N}}.$
Finally, Assumption A2 states that the depth-$L$ planner returns an $\mathcal{O}(1/L)$-optimal value estimate, hence there exists $C_{2}>0$ (scaling linearly with $H$) such that
$R_{\mathrm{pl}}\;\leq\;\frac{C_{2}}{L}.$
Taking a union bound over these three events yields probability at least $1-\delta$ that all bounds hold simultaneously. Combining them and absorbing constants into the big-$\mathcal{O}$ notation gives the claimed result. $\square$
It is worth noting that the regret term $\mathcal{O}\!\bigl(\sqrt{HSA\ln(1/\delta)}\bigr)$ in Theorem 2 is stated in the finite-state, finite-action setting, following the posterior-sampling analysis of [russo2014learning]. In our experiments, however, the state space is continuous and the value functions are represented with function approximators (Gaussian processes and neural networks). A fully rigorous extension of this regret bound to the continuous case would require additional regularity assumptions and technical development, which lies beyond the scope of this work. Accordingly, we view Theorem 2 as an idealized finite-state proxy that captures the qualitative dependence of regret on $H$, NN, and $L$, rather than as a sharp bound for the specific continuous domains studied here.